OpenStudy (anonymous):

Use properties of logs to simplify the ex- pression log4(x −square root (x2 − 28 )) + log4(x + square root (x2 − 28 )). So the property logs for addition is log base b (xy) so I did log base 4 (x-square root of (x^2-28))(logbase4(x+square root of (x^2-28)) Am I supposed to foil this out? Is there a better way of understanding this problem?

OpenStudy (anonymous):

\[\log_{n}x+\log_{n}x=2\log_{n}x \]

OpenStudy (anonymous):

\[=\log_{n}x^2\]

OpenStudy (anonymous):

both logs given are different...I am just confused

OpenStudy (anonymous):

It's difference between two squares.

OpenStudy (anonymous):

\[=\log_{4} (x^2-(x^2-28))\]

OpenStudy (anonymous):

\[=\log_{4}28\]

OpenStudy (anonymous):

the expression is= ln4(x-sqrt(x^2-28))+ln4(x+sqrt(x^2-28)) =ln4[{x-sqrt(x^2-28)}{x+sqrt(x^2-28)}] =ln4[x^2-(x^2-28)] using (a+b)(a-b)=a^2-b^2 =ln4[28] =ln4[(4)(7)] =ln4[4]+ln4[7] =1+ln4[7] this is simplified.

OpenStudy (anonymous):

thank you thank you thank you so much for helping me out!